OpenStudy (angelina__nievse):

Please check my work equivalent to r^m divided by r^n I think r^m + n Question 2. (4^4) ^X = 4^32 what is the value of x I say 8

jimthompson5910 (jim_thompson5910):

`equivalent to r^m divided by r^n I think r^m + n` incorrect

jimthompson5910 (jim_thompson5910):

if it was r^m times r^n, then it would be r^(m+n)

OpenStudy (angelina__nievse):

Would it be r^m-n

jimthompson5910 (jim_thompson5910):

use parenthesis to indicate that m-n is in the exponent r^m-n is completely different than r^(m-n)

jimthompson5910 (jim_thompson5910):

\[\Large r\wedge m - n = r^m - n\] \[\Large r\wedge (m - n) = r^{m - n}\]

jimthompson5910 (jim_thompson5910):

So the answer to the first question is `r^(m-n)`

jimthompson5910 (jim_thompson5910):

as for your second question `(4^4) ^X = 4^32 what is the value of x I say 8` It is correct. x = 8 is the final answer

OpenStudy (angelina__nievse):

Thanks

jimthompson5910 (jim_thompson5910):

no problem